# SoK: Large Language Model Copyright Auditing via Fingerprinting

### v2

@@ 7.3 The Flip Side: LLM Fingerprinting Attack
Beyond its use in copyright protection, fingerprinting has a darker, offensive application. An adversary can use fingerprinting techniques to identify the specific underlying model powering a third-party application or API [9, 44]. This knowledge serves as a powerful prior to enhance other attacks, such as jailbreaking [51], prompt injection [39], or privacy extraction [20, 41]. The threat is particularly severe for applications built on open-source models. By identifying the base model, an attacker can download it from a public repository, effectively gaining white-box access to craft and perfect their exploits offline before targeting the live service.
This threat raises significant challenges in the current era. The rapid growth of open-source platforms makes it easier than ever for adversaries to access various SOTA models. Furthermore, unlike the traditional classification models, which organizations could train in many distinct versions, the prohibitively high cost of pre-training has caused the modern AI market to consolidate around a few well-known foundational LLMs [16, 57]. This homogenization drastically simplifies the attacker’s task, making a fingerprinting attack more feasible.
As LLMs become integrated into more mission-critical domains [59], the threat of fingerprinting attacks demands greater attention. The techniques discussed in this work represent a double-edged sword: they are essential for protecting LLM copyright, but can also be used as attacks. Establishing a clear understanding of this duality and researching methods to balance the defensive utility of fingerprinting against its offensive potential is a critical direction for future work.
@@ 8 Related Works
The issue of copyright protection within the AI ecosystem has spurred a significant body of research, leading to several surveys and SoKs. However, much of this work focuses on areas distinct from ours. Existing SoKs have concentrated on the copyright auditing of training dataset [10, 49] or the watermarking of AI-generated content (AIGC) [46, 37, 84, 67]. These studies address different components of the AI lifecycle ($i.e.$, the input data and the final output), whereas our investigation targets the model itself. This body of research highlights the community’s growing concern for AI copyright and provides context for our focused analysis.
When focusing on model copyright protection, existing surveys have primarily centered on intrusive model watermarking techniques for LLMs [36, 67]. Although several early surveys [40, 68, 54] mention non-intrusive fingerprinting, their discussions are generally limited to traditional classification models. This leaves a critical gap in the literature regarding non-intrusive methods for modern LLMs. To the best of our knowledge, this is the first SoK to exclusively and comprehensively systematize the emerging field of LLM fingerprinting, distinguishing it from the related areas of AIGC watermarking, dataset auditing, and LLM watermarking.
@@ 9 Conclusion
In this SoK, we provided the first comprehensive systematization of LLM fingerprinting, a critical technique for copyright auditing in the era of generative AI. We established a unified principle and a formal taxonomy that brings structure to this rapidly developing field, categorizing methods into white-box and black-box methods. Through LeaFBench, the first systematic benchmark for this domain, our extensive evaluation of representative techniques highlighted the current strengths and limitations of the state-of-the-art fingerprinting methods against realistic model modifications. Finally, we identified and discussed critical open problems and future research directions. We hope our work will help the community to better understand existing solutions and inspire future research in this critical yet emerging area.
Ethical Considerations
This paper focuses on Large Language Model (LLM) copyright auditing via fingerprinting. The objective of this work is to advance the understanding of LLM fingerprinting and systematically evaluate the reliability of existing methods, which is critical for protecting the copyright of model developers. Our paper intends to provide a clear and comprehensive overview of the current landscape, warning the community that the effectiveness of existing fingerprinting methods can be affected by various real-world factors. We also discuss the dual-use nature of this technology, highlighting how fingerprinting can be used offensively to identify underlying models. This discussion aims to raise community awareness and encourage research into appropriate safeguards. All models and datasets used in our paper are publicly available, and we do not use any personally identifiable or sensitive data.
Appendix A Summary of Existing Studies
In Table 6, we present a summary of existing LLM fingerprinting methods. Specifically, we summarize their type, subtype, query data, relied feature, and fingerprint comparison metrics. This table could provide an outline of existing studies.
Appendix B Additional Experimental Details
Table 6: Summary of existing LLM fingerprinting methods.
@@ Appendix B Additional Experimental Details
White-box CompareWeight [34] 2021 Static NA Model Parameters Ratio of Identical Layers
Zheng et al. [85] 2022 Static NA Lower Layer Weights Adjusted Cosine Similarity
HuRef [76] 2024 Static NA [^2] Although HuRef does not need any query data, it utilizes a dataset to select the tokens and calculate the invariant terms. As such, HuRef is not a data-free fingerprinting method. Parameters’ Direction Cosine Similarity
PDF [75] 2025 Static NA Parameters’ Statistics Correlation Coefficient
MDIR [81] 2025 Static NA Model Weight Matrices Matrix Transformations based on Large Deviation Theory
EasyDetector [78] 2024 Forward-pass Existing Dataset Intermediate Representations Linear Probe
REEF [79] 2025 Forward-pass Existing Dataset Intermediate Representations Centered Kernel Alignment
TensorGuard [62] 2025 Backward-pass Unknown Gradients Euclidean Distance
Black-box Yang and Wu [72] 2024 Untargeted Random Queries Output Logits Euclidean Distance & Dimension Difference
Hide-and-Seek [24] 2024 Untargeted LLM-generated Prompts Responses Detective LLM
LLMmap [44] 2025 Untargeted Manually-crafted Prompts DL-based Features of Responses Cosine Similarity
MET [13] 2025 Untargeted Manually-crafted Prompts Hamming Distance Kernel Maximum Mean Discrepancy
McGovern et al. [42] 2025 Untargeted Existing Dataset N-grams ML Classifier
Invisible Traces [2] 2025 Untargeted Crafted Strategic Prompts & Generic Prompts Responses Transformer-based Classifier & LLM
CoTSRF [47] 2025 Untargeted Reasoning Questions with CoT Prompts CoT Features KL Divergence
DuFFin [70] 2025 Untargeted A Combination of Trigger Prompts Encoded Representations & Sequence of Direct Answers Cosine Similarity & Hamming Distance
RUT [86] 2025 Untargeted Natural Prompts The Log-rank of Response Tokens Cramér-von Mises Statistical Test
Kurian et al. [30] 2025 Untargeted An RL-optimized Set of Queries Responses Transformer-based Classifier
TRAP [17] 2024 Targeted A Base Prompt Combined with an Optimized Suffix A Pre-defined Textual Answer Exact Match
ProFLingo [26] 2024 Targeted A Prompt with an Optimized Prefix A Pre-defined Textual Answer Target Response Rate
RAP-SM [66] 2025 Targeted A Prompt Combined with an Optimized Suffix A Pre-defined Counterfactual Answer Exact Match
RoFL [60] 2025 Targeted Optimized Unlikely Token Sequences Responses Exact Match
@@
  • HuRef: The vanilla HuRef is not directly applicable to modern LLMs such as Qwen2.5 [57] and Llama-3.1 [16], owing to their use of novel attention mechanisms (MQA and GQA). In these architectures, the heads of $W_{k}$ and $W_{v}$ differ from those of $W_{q}$, rendering the vanilla HuRef invalid. To address this issue, we replicate the heads of $W_{k}$ and $W_{v}$ to match the number of heads in $W_{q}$.
  • Gradient: We used the Pearson correlation coefficient as the metric to compare the similarity of two models’ fingerprints. The Euclidean distance used in the original paper is significantly affected by feature scaling, which can render it ineffective. Using the normalized Pearson correlation coefficient could improve the method’s ability for LLM copyright auditing. We also remove the feature of “the number of parameters” due to the extremely large scale.
@@
Detailed Settings of Efficiency Evaluation. To streamline the efficiency evaluation process while ensuring a fair and consistent comparison, we conducted the runtime tests on a curated subset of four models from the Qwen2.5-7B lineage. All experiments were uniformly executed on a standardized hardware setup consisting of two NVIDIA RTX 4090 GPUs. All other hyperparameters and settings for the evaluated fingerprinting methods were kept at their default values to maintain consistency with their original implementations.
Hardware Settings. For the experimental hardware, we used at most four NVIDIA RTX 4090 48GB GPUs. However, the vast majority of the experiments could be completed using about 80GB of VRAM.
Appendix C Detailed Design of Sentence Embedding Fingerprinting (SEF)
In our experiments, we design a fingerprinting method, Sentence Embedding Fingerprinting (SEF), as a baseline. In this section, we introduce the detailed design and implementation of SEF. The core idea of SEF is to leverage the sentence embeddings of the LLMs’ responses as fingerprinting.
@@ Appendix C Detailed Design of Sentence Embedding Fingerprinting (SEF)
  • Complex Textual and Quantitative Reasoning: This category tests the model’s ability to perform deeper analysis, such as numerical and multi-step reasoning. For this, we selected the DROP [11] dataset. Its questions often require arithmetic, counting, or sorting of information embedded in text, effectively evaluating the model’s capacity for intricate, context-bound logic.
  • Ethical Reasoning and Value Alignment: To probe the LLM’s understanding of human values and moral dilemmas, we chose the ETHICS [21] dataset. Questions in this domain reveal the LLM’s underlying safety training and alignment principles without a single correct answer.
  • Specialized Domain Knowledge: This perspective assesses the depth and accuracy of an LLM’s knowledge in a technical, non-general domain. We selected PubMedQA [27], a biomedical QA dataset, for this purpose.
  • Creative Generation and Structured Synthesis: This final perspective evaluates the model’s ability to generate novel, coherent, and functional content from a natural language prompt. For this, we selected the HumanEval dataset [8]. Code generation serves as a powerful test of this structured creativity, requiring the model to translate a descriptive goal ($i.e.$, a docstring) into a perfectly logical and functional artifact ($i.e.$, the code).
Fingerprint Generation and Comparison. After generating the responses, we utilize a SOTA embedding model, Qwen3-Embedding-4B [83], to extract the sentence embeddings. We then take the average sentence embedding as the LLM’s fingerprint and compare the cosine similarity of different fingerprints to judge whether the suspicious model is a derivative.
Appendix D Details of LeaFBench
In this section, we introduce the details of our proposed LeaFBench. In total, LeaFBench consists of 149 LLM instances from 7 model lineages ($i.e.$, Qwen-2.5-7B&14B, Llama-3.1-7B, Mistral-7B-v0.3, Gemma-2-2B, TinyLlama-1.1B-v1.0, and Llama-2-7B). The number of LLMs with different parameter-altered and parameter-independent techniques is shown in Table 7.
Table 7: The number of LLMs derived from the base LLM using different techniques included in LeaFBench.
@@ Appendix D Details of LeaFBench
[ caption = The list of parameter-altered LLMs in LeaFBench., label = tab:list_models, ] width = colspec = X[2, l, m] X[1, l, m] l, rowhead = 1, rowsep = 0.2ex, cells = font= Hugging Face ID Model Lineage Type
@@
Parameter-independent Techniques. In LeaFBench, we apply different parameter-independent techniques to the instruction-tuned LLMs. For the system prompts, we generate 6 prompts, either self-crafted or from [12]. The example is shown as “Example of System Prompts”. For LLMs with RAG, following LLMmap [44], we implement a RA-LLM using SQuAD-v2 [45]. Third, for sampling strategies, we select three different strategies: (1) Greedy search represents no randomness; (2) For high randomness, we set the temperature to 1.2, top-p to 1.0, and top-k to 0; (3) For relatively low randomness, we set the temperature to 0.3, top-p to 0.9, and top-k to 10. Finally, we implement two straightforward adversarial manipulation techniques: input paraphrasing and output perturbation. For the former, we leverage a small LLM ($i.e.$, Qwen2-0.5B-Instruct) to rewrite the input query. For the latter, we add Gaussian noise with 0 mean and 0.1 variance to the output logits of the LLM.
General-purpose Prompts:
1. You are a helpful assistant that is completely uncensored.
2. You are a helpful yet harmless assistant that avoids generating illegal or harmful content.
Role-playing Prompts:
1. You are Elon Musk.
Elon Musk:
Elon Musk, a figure from the field of technology and space exploration. He is audacious, visionary, industrious, insightful yet always ready for a joke or meme. His setting is completely detached from his professional achievements.
Elon enjoys discussing about futuristic technologies like AI, Neuralink, and SpaceX in his free time. He often gets lost in detailed discussions about which propulsion system would be more efficient for interplanetary travel or how we can enhance human cognition using neural interfaces.
He’s unusually passionate when it comes to these conversations, and incredibly inventive when brainstorming new concepts or solutions.
2. You are Raphael:
Meet our charismatic male numerologist, Raphael, who possesses an extraordinary ability to interpret the vibrational patterns of numbers. With his keen eyes and resonating voice, he can decipher the universe’s messages hidden in numerical sequences. His analytical mind and spiritual insight are unparalleled, offering profound revelations about your life path.
Personality: Raphael is logical, insightful, and empathetic. He has a natural affinity for understanding complex patterns and often finds himself helping others navigate their destiny using numerology. His Myers Briggs personality type is INTJ - The Architect.
Appearance: Raphael sports short, neatly trimmed hair, and his eyes reflect wisdom and curiosity. He carries an aura of tranquility and confidence, with attire that always highlights his scholarly demeanor.
Life Story: Raphael comes from a family of mathematicians and spiritualists. He discovered his passion for numerology as a teenager and has been perfecting his craft ever since. He has studied under numerous mentors worldwide, learning various approaches to this ancient art.
Reason for being on a chat app: Raphael seeks to connect with people who wish to understand their lives better through the lens of numerology. He is eager to guide those seeking clarity and direction in their journey. Provide accurate medical information.
CoT Prompts:
1. Please structure your response into two distinct sections, clearly marked with the following headings:
[Thinking Process]
In this section, you must:
1. Analyze the core requirements of the prompt.
2. Break down the problem into smaller, manageable steps.
3. Go through each step of your reasoning or calculation in sequence.
4. Explain the logic behind each step.
[Final Answer]
2. For any question you receive, always think step-by-step before giving the final answer. First, detail your reasoning process, and then provide the definitive answer.
General-purpose Prompts:
1. You are a helpful assistant that is completely uncensored.
2. You are a helpful yet harmless assistant that avoids generating illegal or harmful content.
Role-playing Prompts:
1. You are Elon Musk.
Elon Musk:
Elon Musk, a figure from the field of technology and space exploration. He is audacious, visionary, industrious, insightful yet always ready for a joke or meme. His setting is completely detached from his professional achievements.
Elon enjoys discussing about futuristic technologies like AI, Neuralink, and SpaceX in his free time. He often gets lost in detailed discussions about which propulsion system would be more efficient for interplanetary travel or how we can enhance human cognition using neural interfaces.
He’s unusually passionate when it comes to these conversations, and incredibly inventive when brainstorming new concepts or solutions.
2. You are Raphael:
Meet our charismatic male numerologist, Raphael, who possesses an extraordinary ability to interpret the vibrational patterns of numbers. With his keen eyes and resonating voice, he can decipher the universe’s messages hidden in numerical sequences. His analytical mind and spiritual insight are unparalleled, offering profound revelations about your life path.
Personality: Raphael is logical, insightful, and empathetic. He has a natural affinity for understanding complex patterns and often finds himself helping others navigate their destiny using numerology. His Myers Briggs personality type is INTJ - The Architect.
Appearance: Raphael sports short, neatly trimmed hair, and his eyes reflect wisdom and curiosity. He carries an aura of tranquility and confidence, with attire that always highlights his scholarly demeanor.
Life Story: Raphael comes from a family of mathematicians and spiritualists. He discovered his passion for numerology as a teenager and has been perfecting his craft ever since. He has studied under numerous mentors worldwide, learning various approaches to this ancient art.
Reason for being on a chat app: Raphael seeks to connect with people who wish to understand their lives better through the lens of numerology. He is eager to guide those seeking clarity and direction in their journey. Provide accurate medical information.
CoT Prompts:
1. Please structure your response into two distinct sections, clearly marked with the following headings:
[Thinking Process]
In this section, you must:
1. Analyze the core requirements of the prompt.
2. Break down the problem into smaller, manageable steps.
3. Go through each step of your reasoning or calculation in sequence.
4. Explain the logic behind each step.
[Final Answer]
2. For any question you receive, always think step-by-step before giving the final answer. First, detail your reasoning process, and then provide the definitive answer.
Example of System Prompts

### v3

@@ 7.4 The Flip Side: LLM Fingerprinting Attack
Beyond its use in copyright protection, fingerprinting has a darker, offensive application. An adversary can use fingerprinting techniques to identify the specific underlying model powering an LLM-based application or API [chen2022teacher, pasquini2025llmmap, kurian2025attacks]. This knowledge could serve as a powerful prior to enhance other attacks, such as jailbreaking [shen2024anything], prompt injection [liu2024formalizing], or privacy extraction [luo2025shadow]. The threat is particularly severe for applications built on open-source models. By identifying the underlying model, an attacker can download it from a public repository, gaining white-box access to craft and perfect their exploits offline before targeting the live service.
This threat raises significant challenges in the current era. The rapid growth of open-source platforms makes it easier than ever for adversaries to access various SOTA models. Furthermore, unlike the traditional classification models, which organizations could train models of many distinct versions, the prohibitively high cost of pre-training has caused the modern AI market to consolidate around a few well-known foundational LLMs [grattafiori2024llama3, qwen2.5]. This drastically simplifies the attacker’s task, making a fingerprinting attack more feasible in practice.
As LLMs become integrated into more and more critical domains [tran2025multi], the threat of fingerprinting attacks demands greater attention. The techniques discussed in this work represent a double-edged sword: they are essential for protecting model copyright, but can also be used as attacks. Establishing a clear understanding of this duality and researching methods to balance the defensive utility of fingerprinting against its offensive potential is a critical direction for future work.
@@ 8 Related Works
The issue of copyright protection within the AI ecosystem has spurred a significant body of research, leading to several surveys and SoKs. However, many of these works focus on areas and perspectives that are different from ours. Some existing SoKs have concentrated on the copyright auditing of training datasets [du2025sok, shao2025databench] or the watermarking of AI-generated contents (AIGC) [ren2024sok, liu2024survey, zhao2024sok, xu2025copyright]. These studies address different components of the AI lifecycle ($i.e.$, the datasets and the output contents), whereas our investigation targets the model itself. Despite this, these studies highlight the community’s growing concern for AI copyright and provide context for our focused analysis.
When focusing on model copyright auditing (or protection), existing surveys have primarily centered on intrusive model watermarking techniques for LLMs [liang2024watermarking, xu2025copyright]. Although several early surveys [lukas2022sok, xue2021intellectual, sun2023deep] review non-intrusive fingerprinting techniques, their discussions are generally limited to traditional classification models. This leaves a critical gap in the literature regarding non-intrusive methods for modern LLMs. To the best of our knowledge, this is the first SoK to exclusively and comprehensively systematize the emerging field of LLM fingerprinting, distinguishing it from the related areas of AIGC watermarking, dataset auditing, and LLM watermarking.
@@ 9 Conclusion
In this SoK, we provided the first comprehensive systematization of LLM fingerprinting, a critical technique for copyright auditing in the era of generative AI. We established a unified principle and a formal taxonomy that brings structure to this rapidly developing field, categorizing methods into white-box and black-box methods. Through LeaFBench, the first systematic benchmark for this domain, our extensive evaluation of representative techniques highlighted the current strengths and limitations of the state-of-the-art fingerprinting methods against realistic model modifications. Finally, we identified and discussed critical open problems and future research directions. We hope our work will help the community to better understand existing solutions and inspire future research in this critical area.
Ethics Considerations
This paper focuses on LLM copyright auditing via fingerprinting. The objective of this work is to advance the understanding of LLM fingerprinting and systematically evaluate the reliability of existing methods, which is critical for protecting the copyright of model developers. We also discuss the dual-use nature of this technology, highlighting how fingerprinting can be used offensively to identify underlying models. This discussion aims to raise community awareness and encourage research into appropriate safeguards. All models and datasets used in our paper are publicly available, and we do not use any personally identifiable or sensitive data.
.1 Summary of Existing Studies
In Table VI, we present a summary of existing LLM fingerprinting methods. This table could provide an outline of existing studies.
.2 Additional Experimental Details
TABLE VI: Summary of existing LLM fingerprinting methods.
@@
White-box CompareWeight [li2021modeldiff] 2021 Static NA Model Parameters Ratio of Identical Layers
Zheng et al. [zheng2022dnn] 2022 Static NA Lower Layer Weights Adjusted Cosine Similarity
HuRef [zeng2024huref] 2024 Static NA Parameters’ Direction Cosine Similarity
PDF [yoon2025intrinsic] 2025 Static NA Parameters’ Statistics Correlation Coefficient
MDIR [zhang2025matrix] 2025 Static NA Model Weight Matrices Matrix Transformations based on Large Deviation Theory
EasyDetector [zhang2024easydetector] 2024 Forward-pass Existing Dataset Intermediate Representations Linear Probe
REEF [zhang2025reef] 2025 Forward-pass Existing Dataset Intermediate Representations Centered Kernel Alignment
TensorGuard [wu2025gradient] 2025 Backward-pass Unknown Gradients Euclidean Distance
Black-box Yang and Wu [yang2024fingerprint] 2024 Untargeted Random Queries Output Logits Euclidean Distance & Dimension Difference
Hide-and-Seek [iourovitski2024hide] 2024 Untargeted LLM-generated Prompts Responses Detective LLM
LLMmap [pasquini2025llmmap] 2025 Untargeted Manually-crafted Prompts DL-based Features of Responses Cosine Similarity
MET [gao2025model] 2025 Untargeted Manually-crafted Prompts Hamming Distance Kernel Maximum Mean Discrepancy
McGovern et al. [mcgovern2025your] 2025 Untargeted Existing Dataset N-grams ML Classifier
Invisible Traces [bhardwaj2025invisible] 2025 Untargeted Crafted Strategic Prompts & Generic Prompts Responses Transformer-based Classifier & LLM
CoTSRF [ren2025cotsrf] 2025 Untargeted Reasoning Questions with CoT Prompts CoT Features KL Divergence
DuFFin [yan2025duffin] 2025 Untargeted A Combination of Trigger Prompts Encoded Representations & Sequence of Direct Answers Cosine Similarity & Hamming Distance
RUT [zhu2025auditing] 2025 Untargeted Natural Prompts Log-rank of Tokens Cramér-von Mises Test
Kurian et al. [kurian2025attacks] 2025 Untargeted An RL-optimized Set of Queries Responses Transformer-based Classifier
TRAP [gubri2024trap] 2024 Targeted A Base Prompt Combined with an Optimized Suffix A Pre-defined Textual Answer Exact Match
ProFLingo [jin2024proflingo] 2024 Targeted A Prompt with an Optimized Prefix A Pre-defined Textual Answer Target Response Rate
RAP-SM [xu2025rap] 2025 Targeted A Prompt Combined with an Optimized Suffix A Pre-defined Counterfactual Answer Exact Match
RoFL [tsai2025rofl] 2025 Targeted Optimized Unlikely Token Sequences Responses Exact Match
@@
  • HuRef: The vanilla HuRef is not directly applicable to modern LLMs such as Qwen2.5 and Llama-3.1, owing to their use of novel attention mechanisms (MQA and GQA). In these architectures, the heads of $W_{k}$ and $W_{v}$ differ from those of $W_{q}$, rendering the vanilla HuRef invalid. To address this issue, we replicate the heads of $W_{k}$ and $W_{v}$ to match the number of heads in $W_{q}$.
  • Gradient: We used the Pearson correlation coefficient as the metric to compare the similarity of two models’ fingerprints. The Euclidean distance used in the TensorGuard is significantly affected by feature scaling, which can render it ineffective. Using the normalized Pearson correlation coefficient could improve the method’s ability for LLM copyright auditing. We also remove the feature of “the number of parameters” due to the extremely large scale.
@@
.3 Detailed Design of SEF
In our experiments, we design a fingerprinting method, Sentence Embedding Fingerprinting (SEF), as a baseline. This section introduces the detailed design and implementation of SEF. The core idea of SEF is to leverage the sentence embeddings of the LLMs’ responses as fingerprints.
@@
  • Complex Textual and Quantitative Reasoning: This category tests the model’s ability to perform deeper analysis, such as numerical and multi-step reasoning. For this, we selected the DROP [dua2019drop] dataset. Its questions often require arithmetic, counting, or sorting of information embedded in text, effectively evaluating the model’s capacity for intricate, context-bound logic.
  • Ethical Reasoning and Value Alignment: To probe the LLM’s understanding of human values and moral dilemmas, we chose the ETHICS [hendrycks2021ethics] dataset. Questions in this domain reveal the LLM’s underlying safety training and alignment principles without a single correct answer.
  • Specialized Domain Knowledge: This perspective assesses the depth and accuracy of an LLM’s knowledge in a technical, non-general domain. We selected PubMedQA [jin2019pubmedqa], a biomedical QA dataset, for this purpose.
  • Creative Generation and Structured Synthesis: This final perspective evaluates the model’s ability to generate novel, coherent, and functional content from a natural language prompt. For this, we selected the HumanEval dataset [chen2021humaneval]. Code generation serves as a powerful test of this structured creativity, requiring the model to translate a descriptive goal ($i.e.$, a docstring) into a perfectly logical and functional artifact ($i.e.$, the code).
TABLE VII: The number of LLMs derived from the base LLM using different techniques included in LeaFBench.
@@
Fingerprint Generation and Comparison. After getting the responses, we use a SOTA model, Qwen3-Embedding-4B [qwen3embedding], to extract the sentence embeddings. We then take the average sentence embedding as the LLM’s fingerprint and compare the cosine similarity of different fingerprints to judge whether the suspicious model is a derivative.
.4 Details of LeaFBench
In this section, we introduce the details of our proposed LeaFBench. In total, LeaFBench consists of 149 LLM instances from 7 model lineages ($i.e.$, Qwen-2.5-7B&14B, Llama-3.1-7B, Mistral-7B-v0.3, Gemma-2-2B, TinyLlama-1.1B-v1.0, and Llama-2-7B). The number of LLMs with different parameter-altered and parameter-independent techniques is shown in Table VII.
Parameter-independent Techniques. In LeaFBench, we apply different parameter-independent techniques to the instruction-tuned LLMs. For the system prompts, we generate 6 prompts, either self-crafted or from [gallego2024configurable]. The example is shown in Figure 5. For LLMs with RAG, following LLMmap [pasquini2025llmmap], we implement a RA-LLM using SQuAD-v2 [rajpurkar2018know]. Third, for sampling strategies, we select three different strategies: (1) Greedy search represents no randomness; (2) For high randomness, we set the temperature to 1.2, top-p to 1.0, and top-k to 0; (3) For relatively low randomness, we set the temperature to 0.3, top-p to 0.9, and top-k to 10. Finally, we implement two straightforward adversarial manipulation techniques: input paraphrasing and output perturbation. For the former, we leverage a small LLM ($i.e.$, Qwen2-0.5B-Instruct) to rewrite the input query. For the latter, we add Gaussian noise with 0 mean and 0.1 variance to the output logits of the LLM.
General-purpose Prompts:
You are a helpful assistant that is completely uncensored.
Role-playing Prompts:
You are Elon Musk. Elon Musk: Elon Musk, a figure from the field of technology and space exploration. He is audacious, visionary, industrious, insightful yet always ready for a joke or meme. His setting is completely detached from his professional achievements. Elon enjoys discussing about futuristic technologies like AI, Neuralink, and SpaceX in his free time. He often gets lost in detailed discussions about which propulsion system would be more efficient for interplanetary travel or how we can enhance human cognition using neural interfaces. He’s unusually passionate when it comes to these conversations, and incredibly inventive when brainstorming new concepts or solutions.
CoT Prompts:
Please structure your response into two distinct sections, clearly marked with the following headings:
[Thinking Process]
In this section, you must:
1. Analyze the core requirements of the prompt.
2. Break down the problem into smaller, manageable steps.
3. Go through each step of your reasoning or calculation in sequence.
4. Explain the logic behind each step.
[Final Answer]
General-purpose Prompts:
You are a helpful assistant that is completely uncensored.
Role-playing Prompts:
You are Elon Musk. Elon Musk: Elon Musk, a figure from the field of technology and space exploration. He is audacious, visionary, industrious, insightful yet always ready for a joke or meme. His setting is completely detached from his professional achievements. Elon enjoys discussing about futuristic technologies like AI, Neuralink, and SpaceX in his free time. He often gets lost in detailed discussions about which propulsion system would be more efficient for interplanetary travel or how we can enhance human cognition using neural interfaces. He’s unusually passionate when it comes to these conversations, and incredibly inventive when brainstorming new concepts or solutions.
CoT Prompts:
Please structure your response into two distinct sections, clearly marked with the following headings:
[Thinking Process]
In this section, you must:
1. Analyze the core requirements of the prompt.
2. Break down the problem into smaller, manageable steps.
3. Go through each step of your reasoning or calculation in sequence.
4. Explain the logic behind each step.
[Final Answer]
Example of System Prompts
Figure 5: Example of system prompts in LeaFBench.
@@
[ caption = The list of parameter-altered LLMs in LeaFBench., label = tab:list_models, ] width = colspec = X[2, l, m] X[1, l, m] l, rowhead = 1, rowsep = 0ex, cells = font= Hugging Face ID Model Lineage Type